OpenStudy (anonymous):

f(x)= 3x+2 g(x)=3x-5 Find f/g

OpenStudy (anonymous):

1st of all \(\Huge \color{red}{\frak{Good ~job~on~ur~1st~question!!!!!}}\)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Is it f/g (x)=3x+2/(3x-5) domain x|x cannot = 5/3

OpenStudy (anonymous):

i honestly confuse im trying to figure it out touhg

OpenStudy (anonymous):

so whaat are we trying to solve>??

OpenStudy (anonymous):

The domain

OpenStudy (anonymous):

Of x

OpenStudy (anonymous):

ok well this help

OpenStudy (anonymous):

They are like a function and I know there's a domain

OpenStudy (anonymous):

idk then srry

OpenStudy (anonymous):

It's ok ^_^